OpenStudy (anonymous):

What is the area of a circle with diameter 12 inches? 144pi in2 6pi in2 12pi in2 36pi in2 @Venny

OpenStudy (anonymous):

Okay the area of a circle is given by the formula \(\Large\ A=\pi(r)^2\) but we are given the diameter and \(\large\ d=2*r\) so if d=12, what is r?

OpenStudy (anonymous):

3.14?

OpenStudy (anonymous):

Nooo, divide 12 by 2, 3.14 is pi

OpenStudy (anonymous):

6

OpenStudy (anonymous):

lol srry

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

No, you forgot to plug it into the formula first \(\Large\ A=\pi(6)^2\)

OpenStudy (anonymous):

6x2=12

OpenStudy (anonymous):

37.68

OpenStudy (anonymous):

Hmm not quite \(\large6^2\) means you need to do \(\large6*6\) because you have to multiply the number times itself

OpenStudy (anonymous):

36 is the answer?

OpenStudy (anonymous):

Yes \(36\pi\)